OpenStudy (anonymous):

A truck is at a stoplight. When the light turns green, it accelerates at 2.5m/s^2. At the same instant, a car passes the truck going 15m/s. Where and when does the truck catch up with the car?

OpenStudy (anonymous):

\[a*t^{2}/2=v*t\] t=15*2/2.5=12 seconds

OpenStudy (anonymous):

how do you get a * t^2 /2 ?

OpenStudy (amistre64):

its a common equation of motion

OpenStudy (amistre64):

\[distance=\frac12at^2+v_ot\] when does distance equal distance? \[\frac12a_1t^2+v_{1o}t=\frac12a_2t^2+v_{2o}t\] assuming the car is traveiling at a constant speed, its accelration is zero, and since the truck starts from a stop, its initial velocity is zero \[\frac12a_1t^2=v_{2o}t\] \[\frac12a_1t^2-v_{2o}t=0\] if we need to, we can use the quadratic equation, but that does seem to be a bit of overkill .... but i am a math major soooo; \[t=\frac{v_{2o}\pm\sqrt{(v_{2o})^2-4(\frac12a_1)}}{2(\frac12a_1)}\]

OpenStudy (amistre64):

\[\frac12a_1t^2=v_{2o}t\] \[\frac12a_1t=v_{2o}\] \[t=\frac{2v_{2o}}{a_1}\]